# Sparse Hyperparametric Itakura-Saito NMF via Bi-Level Optimization

### v2

@@ 2. Dynamical system approach on H
An approach to solve the bi-level Problem (5) over $\bm{H}$ is to replace the inner problem with a dynamical system [14, 15, 22] and compute an approximation solution.
@@
The idea of the bi-level optimization strategy is to use the IVP-$\Phi$ to approximate the solution of the Problem (5). We do so by solving the following minimization
$\begin{array}[]{rl}\displaystyle\lambda_{l}^{*}=\operatorname*{\textrm{argmin}\,}_{\lambda_{l}}&\mathcal{r}(\lambda_{l})\\ \,\textrm{s.t.}{}{}&\underline{\bm{h}}_{l}^{t}=\Phi_{t}(\underline{\bm{h}}_{l}^{t-1},\lambda_{l})\quad\text{for }t\in[T]\end{array}$ (6)
@@ 3. Hypergradient
We would like to compute the hypergradient, i.e. the gradient of $\mathcal{R}$ with respect to $\bm{\lambda}$, in order to use a gradient descent approach on $\bm{\lambda}$. The hypergradient $\nabla_{\bm{\lambda}}\mathcal{R}$, by using chain rule, is
@@
It is well known that the computation of the hypergradient can be done using Reverse-Mode Differentiation (RMD) or Forward-Mode Differentiation (FMD). Since RMD requires storing specific variables across all iterations and indices in memory, in this work we use FMD, making it more suitable for scenarios where the total quantity of interest is small. For details we refer the reader to [8, 15].
@@ 4 Derivation of the algorithm SHINBO
Now denote $H_{ij}$ the $(i,j)$-th element of the matrix $\bm{H}$, and recall that $\odot$ is the Hadarmard product, then by the complementary slackness $\Psi_{ij}H_{ij}=0$ in the KKT conditions, we get
@@
$H_{ij}=H_{ij}\Big{[}\bm{W}^{\top}(\bm{W}\bm{H})^{\odot-2}\bm{X}\Big{]}_{ij}\bigg{/}\Big{[}\bm{W}^{\top}(\bm{W}\bm{H})^{\odot-1}+2\textrm{Diag}(\bm{\lambda})^{2}\bm{H}\bm{E}\Big{]}_{ij},$ (11)
@@ 4.1 The implementation of the bi-level approach in SHINBO
    where the derivative is computed with respect to the $(l,j)$-th element of $\bm{H}$, and
@@ 5 Experimental Results
In this section, we present the numerical results of comparing the proposed algorithm SHINBO with the multiplicative update (MU) algorithm [12]. We also compare SHINBO with the penalized MU, which is the update (12) under a fixed penalty hyperparameter $\bm{\lambda}$ with $\lambda_{1}=\lambda_{2}=\ldots=\lambda_{r}$ for every row of $\bm{H}$. We summarize their difference in the table below.
@@
We evaluate the algorithms on two datasets: one generated synthetically, and the other one obtained by processing real vibrational signals measured in a laboratory condition from damage rolling bearing elements.
  • 1.
    The synthetic dataset is generated starting from a full-rank decomposition $\bm{X}\approx\bm{W}\bm{H}$, where the factor matrix $\bm{W}$ contains 10% nonzeros in the $mr$ entries and the factor matrix $\bm{H}$ contains 70% nonzeros in the $rn$ entries. In this dataset, we have $(m,n,r)=(100,7,3)$.
  • 2.
    The vibrational signals were measured on the test rig presented in Fig. 1. The platform is equipped with an electric motor, gearbox, couplings, and two bearings. One of the latter was deliberately damaged. Diagnostic signals were measured with the accelerometer (KISTLER Model 8702B500), stacked horizontally to the shaft bearing. The 40-second-long vibration signal was recorded with the sampling rate of 50 kHz. For easier visualization, one second excerpt was selected, and then transformed to a spectrogram using the 128-window length, 100 sample overlapping, and 512 frequency bins. The raw observed signal and its spectrogram are shown in Fig. 2. Assuming $r=4$ (four main components of the mixed signal), we have $(m,n,r)=(257,1782,4)$.
@@ 5.1 Experimental setup
To ensure a fair comparison, all algorithms were initialized using the same initial factor matrices, which were generated from an unpenalized MU-based NMF warm start, that is initialized via nonnegative double singular value decomposition (NNDSVD) of the matrix $\bm{X}$ [4].
@@
[…]
where $k$ indicates the iteration of the outer loop of the algorithm. For the inner loop of the bi-level approach on iteration counter $t$, we stop at the maximum number of (inner) $T=4$ iterations.
@@
For the synthetic dataset, we evaluate the quality of the factorization with respect to the identification problem for the synthetic dataset using the Signal-to-Interference Ratio (SIR) measure [5] between the estimated signals and the true signals. This is a log-scale similarity measure (expressed in decibels), often used in signal processing applications, and its higher value indicates a higher similarity level. To highlight the effectiveness of the proposed method, we conduct a statistical comparison of SIR values across algorithms using the Kruskal-Wallis test, followed by a post-hoc multiple comparison based on the Mann-Whitney test [21] with Benjamini-Hochberg (BH) correction [2], maintaining a significance level of $\alpha=0.05$. We also investigate the sparsity of the results obtained for the synthetic dataset using the following measure of sparsity for a generic matrix $\bm{A}\in\mathbb{R}^{m\times n}$:
@@
On real-world dataset we check the goodness of the factorization by quantifying the impulsive and cyclic behavior of the signal under analysis, using a modified envelope spectrum based indicator (ENVSI) [17, 3] on time profiles. ENVSI can be expressed as spectrum based indicator (SBI), defined as:
@@
where $\text{AIS}_{i}$ is the magnitude of the $i$-th harmonic of the estimated signal in the frequency domain, $\text{S}_{k}$ is the magnitude in the $k$-th frequency bin in the spectrum of the time profile, $M_{1}$ is the number of harmonics to be analyzed (assuming a periodic signal), and $M_{2}$ is the number of frequency bins to calculate the total energy.
SBI is zero if there are no impulsive components in the time profile, whereas a larger SBI occurs when the impulses in the time profile are stronger (which corresponds to the amplitudes in the spectrum) and the noise is weaker. In the experiments, the number of harmonics $M_{1}$ is set to 6, and it was experimentally found to be sufficient for demonstrating the impulsive and period behavior of the SOI representing time profile in the analyzed application. To demonstrate the superiority of the proposed method, we statistically compare ENVSI values across algorithms using a Kruskall-Wallis test and a post-hoc multiple comparison based on Mann-Whitney test with a BH correction, with a statistical significance level set at $\alpha=0.05$.
@@ 5.2 Results on synthetic dataset
Refers to Figure 3, the proposed algorithm demonstrates superior performance in terms of convergence rate of Response function compared to the other methods. We remark that, despite SHINBO exhibited similar and rapid convergence as unpenalized MU for the first 450 iterations (on average), the obtained decompositions have a different SIR value. See the discussion below.
@@
Refers to Figure 4, we can see that SHINBO obtain the best SIR values on both matrices and the higher sparsity on $\bm{H}$ compared to the other methods. On average, SHINBO achieved 10% better SIR and 5% better sparsity on $\bm{H}$ than other algorithms.
Figure 4: Results of sparsity and SIR on the factor matrices on the synthetic dataset. Left two columns are the sparsity of $\bm{W}$ (left) and $\bm{H}$ (right) of the four algorithms: from the left to right are MU with $\lambda=0$, MU with $\lambda=0.1$, MU with $\lambda=0.5$ and SHINBO. The right two columns are the SIR of $\bm{W}$ (left) and $\bm{H}$ (right) of the four algorithms: from the left to right are MU with $\lambda=0$, MU with $\lambda=0.1$, MU with $\lambda=0.5$ and SHINBO.
These results are also confirmed by the statistical comparisons. The Kruskall-Wallis tests present $p$-values lesser that $10^{-14}$ and the details of the pair-wise comparison with a BH comparison are presented in the tables 2 and 3.
Table 2: $p$-values results of pairwise comparisons for SIR coefficients on ($\bm{W},\,\bm{H}$). In the table $\epsilon=10^{-16}$.
@@
MU $\lambda=0.1$ $8.1\cdot 10^{-5}$, $6.3\cdot 10^{-10}$
MU $\lambda=0.5$ $<2\epsilon$,$<2\epsilon$ $<2\epsilon$,$<2\epsilon$
SHINBO 0.14,0.0042 $5.4\cdot 10^{-5},1.2\cdot 10^{-9}$ $<2\epsilon$, $<2\epsilon$
Table 3: $p$-values results of pairwise comparisons for sparsity coefficients on $(\bm{W},\bm{H})$.
@@
MU $\lambda=0.5$ $<2\epsilon$, $<2\epsilon$ $<2\epsilon$,$<2\epsilon$
SHINBO 0.023, $<2\epsilon$ $3.12\cdot 10^{-6}$, $<2\epsilon$ $<2\epsilon,<2\epsilon$

### v3

@@ 2. Dynamical system approach on H
An approach to solve the bi-level Problem (5) over $\bm{H}$ is to replace the inner problem with a dynamical system [21, 22] and compute an approximation solution.
@@
Bi-level optimization uses the IVP-$\Phi$ to approximate the solution of the Problem (5). We do so by solving the following minimization problem
$\begin{array}[]{rl}\displaystyle\lambda_{l}^{*}=\operatorname*{\textrm{argmin}\,}_{\lambda_{l}}&\mathcal{r}(\lambda_{l})\\ \,\textrm{s.t.}&\underline{\bm{h}}_{l}^{t}=\Phi_{t}(\underline{\bm{h}}_{l}^{t-1},\lambda_{l})\quad\text{for }t\in[T],\end{array}$ (6)
@@ 3. Hypergradient
We would like to compute the hypergradient, i.e., the gradient of $\mathcal{R}$ with respect to $\bm{\lambda}$, in order to use a gradient descent approach on $\bm{\lambda}$. The hypergradient $\nabla_{\bm{\lambda}}\mathcal{R}$, by using chain rule, is
@@
It is well known that the computation of the hypergradient can be done using Reverse-Mode Differentiation (RMD) or Forward-Mode Differentiation (FMD). Since RMD requires storing specific variables across all iterations and indices in memory, in this work, we use FMD, making it more suitable for scenarios where the total quantity of interest is small. For details we refer the reader to [5, 21].
@@ 4 Derivation of the algorithm SHINBO
Now denote $H_{ij}$ the $(i,j)$th element of the matrix $\bm{H}$, and recall that $\odot$ is the Hadarmard product, then by the complementary slackness $\Psi_{ij}H_{ij}=0$ in the KKT conditions, we get
@@
$H_{ij}=H_{ij}\Big[\bm{W}^{\top}(\bm{W}\bm{H})^{\odot-2}\bm{X}\Big]_{ij}\bigg/\penalty 50\Big[\bm{W}^{\top}(\bm{W}\bm{H})^{\odot-1}+2\textrm{Diag}(\bm{\lambda})^{2}\bm{H}\bm{E}\Big]_{ij},$ (11)
@@ 4.1 The implementation of the bi-level approach in SHINBO
    where the derivative is computed with respect to the $(l,j)$th element of $\bm{H}$, and
@@
Computational cost
The computational cost of the proposed algorithm depends on the matrix-vector multiplications, as is usual in NMF algorithms, and on the bi-level framework, which requires the use of internal loops. The total computational cost of SHINBO is $\mathcal{O}(KTmnr)$, where $K$ is the number of iterations, and $T$ is the number of internal iterations. Although the proposed method appears to suffer from a high computational cost in construction, the gain in efficiency and accuracy in hyperparameter estimation justifies the high cost. Indeed, solving the NMF problem by running a well-known grid or random search methods requires more effort than SHINBO, which automatically selects the best hyperparameter vector for each row of $\bm{H}$.
@@ 5 Experimental Results
In this section, we present the numerical results of comparing the proposed algorithm SHINBO with the multiplicative update (MU) algorithm in [18]. We also compare SHINBO with the penalized MU, which is the update (12) under a fixed penalty hyperparameter $\bm{\lambda}$ with $\lambda_{1}=\lambda_{2}=\ldots=\lambda_{r}$ for every row of $\bm{H}$. We summarize their difference in the table below.
@@
We evaluate the algorithms on two datasets: one synthetically generated and the other obtained from real vibrational signals measured under laboratory conditions from damaged rolling bearing elements.
  • The synthetic dataset is generated starting from a full-rank decomposition $\bm{X}\approx\bm{W}\bm{H}$, where the factor matrix $\bm{W}$ contains 10% nonzeros in the $mr$ entries and the factor matrix $\bm{H}$ contains 70% nonzeros in the $rn$ entries. In this dataset, we have $(m,n,r)=(100,70,3)$.
  • The vibrational signals were measured on the test rig presented in Fig. 1 (taken from the previous study [14]). The platform is equipped with an electric motor, gearbox, couplings, and two bearings. One of the latter was deliberately damaged. Diagnostic signals were measured with the accelerometer (KISTLER Model 8702B500), stacked horizontally to the shaft bearing. The 40-second-long vibration signal was recorded with a sampling rate of 50 kHz. For easier visualization, one-second excerpt was selected and then transformed to a spectrogram using a 128-window length, 100-sample overlapping, and 512 frequency bins. The raw observed signal and its spectrogram are shown in Fig. 2. Assuming $r=4$ (four main components of the mixed signal), we have $(m,n,r)=(257,1782,4)$.
@@ 5.1 Experimental setup
For synthetic dataset, to ensure a fair comparison, all algorithms were initialized using the same initial factor matrices, which were generated from an unpenalized MU-based NMF warm start, that is initialized via nonnegative double singular value decomposition (NNDSVD) of the matrix $\bm{X}$ [23]. For real-dataset, we initialize the initial factor matrices as follows. Let $|\mathcal{N}|$ denote the truncated Gaussian distribution (i.e., each entries of the matrix follows $\mathcal{N}(0,1)$ and then perform a clipping to replace negative values to 0), we generate both $\bm{W}_{0},\bm{H}_{0}$ using $(1.5|\mathcal{N}|+0.5)/2$. See Fig. 3.
Figure 3: Initial factors generated using $(1.5|\mathcal{N}|+0.5)/2$. Left: $\bm{W}_{0}\bm{H}_{0}$. Right: each rank-1 component in $\bm{W}_{0}\bm{H}_{0}$.
@@
$\big|D_{0}(\bm{X},\bm{W}^{k+1}\bm{H}^{k+1})-D_{0}(\bm{X},\bm{W}^{k}\bm{H}^{k})\big|\,\big/\penalty 50\,\big|D_{0}(\bm{X},\bm{W}^{k}\bm{H}^{k})\big|\,\leq\,10^{-6},$
where $k$ indicates the iteration of the outer loop of the algorithm. For the inner loop of the bi-level approach on iteration counter $t$, we stop at the maximum number of (inner) $T=4$ iterations. We remark that Problem (3) is nonconvex, thus we do not have a theoretical guarantee that the algorithm’s convergence can avoid poor local minima. We only empirically show that the local minima reached are stable across different random datasets.
@@
For the synthetic dataset, we evaluate the quality of the factorization with respect to the identification problem using the Signal-to-Interference Ratio (SIR) measure [24] on both matrices [^2] This measure, computed between the estimated signals and the true signals, is a log-scale similarity measure (expressed in decibels), often used in signal processing applications, and its higher value indicates a higher similarity level.. To highlight the effectiveness of the proposed method, we conduct a statistical comparison of SIR values across algorithms using the Kruskal-Wallis test, followed by a post-hoc multiple comparison based on the Mann-Whitney test [25] with Benjamini-Hochberg (BH) correction [26], maintaining a significance level of $\alpha=0.05$. We also investigate the sparsity of the results obtained for the synthetic dataset using the following measure of sparsity for a generic matrix $\bm{A}\in\mathbb{R}^{m\times n}$:
@@
On the real-world dataset, we check the goodness of the factorization by quantifying the impulsive and cyclic behavior of the signal under analysis, using a modified envelope spectrum-based indicator (ENVSI) [27, 28] on time profiles. ENVSI can be expressed as a spectrum-based indicator (SBI):
@@
where $\text{AIS}_{i}$ is the magnitude of the $i$th harmonic of the estimated signal in the frequency domain, $\text{S}_{k}$ is the magnitude in the $k$th frequency bin in the spectrum of the time profile, $M_{1}$ is the number of harmonics to be analyzed (assuming a periodic signal), and $M_{2}$ is the number of frequency bins to calculate the total energy. SBI is zero if there are no impulsive components in the time profile, whereas a larger SBI occurs when the impulses in the time profile are stronger (which corresponds to the amplitudes in the spectrum), and the noise is weaker. In the experiments, the number of harmonics $M_{1}$ is set to 6, and it was experimentally found to be sufficient for demonstrating the impulsive and periodic behavior of the SOI representing the time profile in the analyzed application. To demonstrate the superiority of the proposed method, we statistically compare ENVSI values across algorithms using a Kruskal-Wallis test and a multiple comparison based on the Mann-Whitney test with BH correction, with a statistical significance level $\alpha=0.05$.
@@ 5.2 Results on synthetic dataset
Referring to Fig.4, the proposed algorithm demonstrates superior performance in terms of the convergence rate of Response function compared to the other methods. We remark that, despite SHINBO exhibiting similar, rapid convergence to unpenalized MU for the first 450 iterations (on average), the resulting decompositions have different SIR values. Referring to Fig.4, we observe that SHINBO, by operating on the vector $\bm{\lambda}$ across its components, achieves the highest SIR values on both matrices, confirming its effectiveness in the identification problem. At the same time, the results show higher sparsity on $\bm{H}$ and lower sparsity on $\bm{W}$, indicating that the automatically selected hyperparameter behaves as expected by enforcing sparsity primarily on the rows of $\bm{H}$. On average, SHINBO outperforms the comparison algorithms with about 10% higher SIR and 5% higher sparsity on $\bm{H}$.
@@
Figure 5: Results of sparsity and SIR on the factor matrices on the synthetic dataset. The left two columns are the sparsity of $\bm{W}$ (left) and $\bm{H}$ (right) of the four algorithms: from left to right are MU with $\lambda=0$, MU with $\lambda=0.1$, MU with $\lambda=0.5$, and SHINBO. The right two columns are SIR of $\bm{W}$ (left) and $\bm{H}$ (right) of the four algorithms: from the left to right are MU with $\lambda=0$, MU with $\lambda=0.1$, MU with $\lambda=0.5$, and SHINBO.
These results are also confirmed by the statistical comparisons. The Kruskal-Wallis tests present $p$-values lesser than $10^{-14}$ and the details of the pairwise comparison with a BH comparison are presented in the tables 2 and 3. Table 2 reports the results of pairwise comparisons for the SIR coefficients on ($\bm{W}$, $\bm{H}$). The significantly low $p$-values indicate that SHINBO consistently achieves higher SIR values compared to MU under different penalization settings, thereby confirming its superior accuracy in handling the identification problem in matrix reconstruction. Similarly, Table 3 presents the $p$-values obtained from pairwise comparisons of sparsity coefficients on ($\bm{W}$, $\bm{H}$). The results show that SHINBO exhibits statistically significant differences with respect to MU variants, highlighting its ability to balance accuracy and sparsity in both synthetic and real datasets.
Table 2: $p$-values of pairwise comparisons for SIR coefficients on ($\bm{W},\,\bm{H}$). In the table $\epsilon=10^{-16}$.
@@
MU $\lambda=0.1$ $8.1\cdot 10^{-5}$, $6.3\cdot 10^{-10}$ - -
MU $\lambda=0.5$ $2\epsilon$$2\epsilon <$$2\epsilon$$2\epsilon -$
SHINBO 0.14,0.0042 $5.4\cdot 10^{-5},1.2\cdot 10^{-9}$ $2\epsilon$, $2\epsilon$
Table 3: $p$-values of pairwise comparisons for sparsity coefficients on $(\bm{W},\bm{H})$.
@@
MU $\lambda=0.5$ $2\epsilon$, $2\epsilon <$$2\epsilon$$2\epsilon -$
SHINBO 0.023, $2\epsilon$ $3.12\cdot 10^{-6}$, $2\epsilon <$$2\epsilon,<2\epsilon$
Sensitivity to the rank
[…]
  • First, MU $\lambda=0.5$ has the smallest variance across all methods, but also at the same time has the lowest SIR on average, meaning the method has the worst performance.
  • The low variance of MU $\lambda=0.5$ is contributed to the fact that the method has a relatively larger $\lambda$ compared to the baseline MU method. Recall that the diversity measure (2), such as $J(\bm{A})=\sum_{i=1}^{n}\|\bm{A}_{i:}\|_{1}^{2}$, is applied on $\bm{H}$, thus a higher $\lambda$ in MU will made the method focusing more on minimizing the column norm of $\bm{H}$. A solution with small column-wise norm will naturally has smaller variance, and thus this explains why MU $\lambda=0.5$ has the smallest variance. In general, if $\lambda$ is huge (e.g. $\lambda=10,100,$...), then we can expect that MU with such value will have even smaller variance on the SIR results. As Problem (3) contains two terms, thus a large $\lambda$ will lead to the MU focusing on column norm minimization instead of the factorization, leading to a lower SIR in reconstruction.
  • Among the MU methods, MU with $\lambda=0$ has the highest SIR on average and also the highest variance. Compared with MU, SHINBO has a higher SIR on average in this case.
[…]
rank MU MU $\lambda=0.1$ MU $\lambda=0.5$ SHINBO
4 $39.096\hskip-2.84526pt\pm\hskip-2.84526pt14.023$ $35.899\hskip-2.84526pt\pm\hskip-2.84526pt11.641$ $17.457\hskip-2.84526pt\pm\hskip-2.84526pt7.164$ $44.734\hskip-2.84526pt\pm\hskip-2.84526pt17.894$
5 $29.830\hskip-2.84526pt\pm\hskip-2.84526pt12.216$ $28.183\hskip-2.84526pt\pm\hskip-2.84526pt11.050$ $14.769\hskip-2.84526pt\pm\hskip-2.84526pt6.545$ $35.801\hskip-2.84526pt\pm\hskip-2.84526pt16.903$
6 $3.423\hskip-2.84526pt\pm\hskip-2.84526pt9.159 22$$2.576\hskip-2.84526pt\pm\hskip-2.84526pt8.694$ $11.849\hskip-2.84526pt\pm\hskip-2.84526pt5.028$ $26.930\hskip-2.84526pt\pm\hskip-2.84526pt13.764$
[…]
rank MU MU $\lambda=0.1$ MU $\lambda=0.5$ SHINBO
4 $41.836\hskip-2.84526pt\pm\hskip-2.84526pt17.689$ $39.183\hskip-2.84526pt\pm\hskip-2.84526pt15.542$ $20.736\hskip-2.84526pt\pm\hskip-2.84526pt10.644$ $46.108\hskip-2.84526pt\pm\hskip-2.84526pt20.023$
5 $30.665\hskip-2.84526pt\pm\hskip-2.84526pt15.257$ $29.489\hskip-2.84526pt\pm\hskip-2.84526pt13.976$ $17.347\hskip-2.84526pt\pm\hskip-2.84526pt9.014$ $36.755\hskip-2.84526pt\pm\hskip-2.84526pt18.479$
6 $23.879\hskip-2.84526pt\pm\hskip-2.84526pt10.719$ $23.317\hskip-2.84526pt\pm\hskip-2.84526pt10.242$ $14.153\hskip-2.84526pt\pm\hskip-2.84526pt6.501$ $27.983\hskip-2.84526pt\pm\hskip-2.84526pt15.461$
Robustness under measurement noise
To assess robustness to measurement noise, we corrupted the clean data $\bm{X}$ by adding i.i.d. zero-mean Gaussian noise and then projecting onto the nonnegative orthant. Formally, we generated
$\bm{Y}=\max(\bm{X}+\varepsilon\mathcal{N}(0,1),0),$
where $\varepsilon>0$ denotes the noise standard deviation. The SIR metrics on both factors in Tables 6-7 show a gradual degradation of SIR as $\varepsilon$ increases (i.e., as the noise level grows), without abrupt performance breakdowns. Across all tested noise conditions, SHINBO remains comparable to the MU baselines, often slightly superior at a low noise level (e.g., $\varepsilon=0.01$), and on par at higher noise, thus substantiating its robustness to measurement noise.
[…]
$\varepsilon$ MU MU $\lambda=0.1$ MU $\lambda=0.5$ SHINBO
0.01 $27.3259\hskip-2.84526pt\pm\hskip-2.84526pt4.7506$ $27.3571\hskip-2.84526pt\pm\hskip-2.84526pt4.5884$ $20.9436\hskip-2.84526pt\pm\hskip-2.84526pt4.4491$ $27.6639\hskip-2.84526pt\pm\hskip-2.84526pt5.1397$
0.05 $19.5527\hskip-2.84526pt\pm\hskip-2.84526pt3.7249$ $19.5402\hskip-2.84526pt\pm\hskip-2.84526pt3.7146$ $16.3236\hskip-2.84526pt\pm\hskip-2.84526pt4.3636$ $19.2795\hskip-2.84526pt\pm\hskip-2.84526pt4.1752$
0.1 $15.6428\hskip-2.84526pt\pm\hskip-2.84526pt3.3477$ $15.6939\hskip-2.84526pt\pm\hskip-2.84526pt3.2599$ $12.6351\hskip-2.84526pt\pm\hskip-2.84526pt4.5464$ $15.3858\hskip-2.84526pt\pm\hskip-2.84526pt3.3559$
[…]
$\varepsilon$ MU MU $\lambda=0.1$ MU $\lambda=0.5$ SHINBO
0.01 $28.3267\hskip-2.84526pt\pm\hskip-2.84526pt4.7552$ $28.4009\hskip-2.84526pt\pm\hskip-2.84526pt4.7298$ $23.8437\hskip-2.84526pt\pm\hskip-2.84526pt6.1478$ $29.0103\hskip-2.84526pt\pm\hskip-2.84526pt4.8775$
0.05 $19.8840\hskip-2.84526pt\pm\hskip-2.84526pt3.6120$ $19.8576\hskip-2.84526pt\pm\hskip-2.84526pt3.6338$ $17.3408\hskip-2.84526pt\pm\hskip-2.84526pt5.4629$ $19.3763\hskip-2.84526pt\pm\hskip-2.84526pt4.1957$
0.1 $15.7853\hskip-2.84526pt\pm\hskip-2.84526pt3.1266$ $15.8039\hskip-2.84526pt\pm\hskip-2.84526pt3.1181$ $13.2717\hskip-2.84526pt\pm\hskip-2.84526pt5.1220$ $15.6706\hskip-2.84526pt\pm\hskip-2.84526pt3.1455$